OpenStudy (anonymous):

Alright, I don't want someone to tell me the answer, just want help to start it? :) 3z^2+12z/z^4

jimthompson5910 (jim_thompson5910):

The problem is \[\Large \frac{3z^2+12z}{z^4}\] right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

what do 3z^2, 12z and z^4 all have in common (in terms of factors)?

OpenStudy (anonymous):

z

jimthompson5910 (jim_thompson5910):

so factor that out of the numerator 3z^2 + 12z to get ???

OpenStudy (anonymous):

3(z^2+4) ?

jimthompson5910 (jim_thompson5910):

factor out z not 3

OpenStudy (anonymous):

i don't know then..

jimthompson5910 (jim_thompson5910):

z^2 is the same as z*z

jimthompson5910 (jim_thompson5910):

so 3z^2 = 3*z*z = z*3z

jimthompson5910 (jim_thompson5910):

12z is z*12

jimthompson5910 (jim_thompson5910):

3z^2 + 12z z*3z + z*12 z(3z + 12) ... using distributive property

OpenStudy (anonymous):

ok, so then what?

jimthompson5910 (jim_thompson5910):

z^4 is the same as z*z^3

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{3z^2+12z}{z^4}\] turns into \[\Large \frac{z(3z + 12)}{z*z^3}\]

jimthompson5910 (jim_thompson5910):

what cancels?

OpenStudy (anonymous):

z cancels out?

OpenStudy (anonymous):

so its 3z+12/z^3 ?

jimthompson5910 (jim_thompson5910):

correct, \[\Large \frac{3z + 12}{z^3}\]

jimthompson5910 (jim_thompson5910):

so when it comes to simplifying things like this, you factor as much as possible and look for any cancellations

OpenStudy (anonymous):

great thank you so much!! :)

jimthompson5910 (jim_thompson5910):

you're welcome